# RL of Thoughts: Navigating LLM Reasoning with Inference-time Reinforcement Learning

### v2

@@ Abstract
Despite rapid advancements in large language models (LLMs), the token-level autoregressive nature constrains their complex reasoning capabilities. To enhance LLM reasoning, inference-time techniques, including Chain/Tree/Graph-of-Thought(s), successfully improve the performance, as they are fairly cost-effective by guiding reasoning through external logical structures without modifying LLMs’ parameters. However, these manually predefined, task-agnostic frameworks are applied uniformly across diverse tasks, lacking adaptability. To improve this, we propose RL-of-Thoughts (RLoT), where we train a lightweight navigator model with reinforcement learning (RL) to generate task-adaptive logical structures at inference time, enhancing LLM reasoning. Specifically, we design five basic logic blocks from the perspective of human cognition. During the reasoning process, the trained RL navigator dynamically selects the suitable logic blocks and combines them into task-specific logical structures according to problem characteristics. Experiments across multiple reasoning benchmarks (AIME, MATH, GPQA, etc.) with multiple LLMs (GPT, Llama, Qwen, and DeepSeek) illustrate that RLoT outperforms established inference-time techniques by up to 13.4%. Remarkably, with less than 3K parameters, our RL navigator is able to make sub-10B LLMs comparable to 100B-scale counterparts. Moreover, the RL navigator demonstrates strong transferability: a model trained on one specific LLM-task pair can effectively generalize to unseen LLMs and tasks. Our code is open-source at https://anonymous.4open.science/r/RL-LLM-Reasoning-1A30.
@@ 3.2 LLM Reasoning as MDP
A2: Clarity for further reasoning
B2: Alternative methods in
further reasoning
C2: Completeness within the step
@@ 3.3 Training of the Navigator Model
Within the MDP framework outlined above, given a specific LLM type and a kind of reasoning task, we train the navigator model. To enhance the learning for challenging reasoning tasks, we extract hard questions from the training set of the target task, i.e., questions that the LLM cannot answer when directly prompted. Then, we use these problems for training the navigator model, from which we randomly select one in each episode and repeat it multiple times. We employ the Double-Dueling-DQN algorithm (Van Hasselt et al., 2016; Wang et al., 2016), and include the online and target network technique to improve the stability during training (Mnih et al., 2015).
@@ 4.1 Experimental Settings
Implementation of RLoT. In our implementation, the navigator model is a simple three-layer multilayer perceptron (MLP) with the Dueling Network architecture (Wang et al., 2016). The model merely contains 2,566 parameters in total, where the lightweight design ensures efficient training and inference. Following existing works like CoT-SC, we perform multiple repeated inferences for each task and filter out the trails that do not meet self-consistency, enhancing the robustness of reasoning.
@@
DeepSeek-R1-
Distill-Qwen-7B
Direct QA 46.67 60.00 92.27 95.74 54.47 83.61 79.48 73.18
@@ 4.2 Overall Performance
We train the navigator model with Qwen2.5-14B-Instruct on the MATH benchmark, where the learning curve shown in Appendix A indicates good convergence. We use the obtained navigator model to enhance the reasoning of multiple LLMs across different reasoning tasks and present the results in Table 2. For the baseline performance, we prioritize using the results reported in the official technical reports of each LLM (Yang et al., 2024a; Dubey et al., 2024; Hurst et al., 2024), if available. Otherwise, we evaluate the performance through our own experiments following the settings established in previous works (Fu et al., 2023; Yang et al., 2024b; Yao et al., 2023).
@@ 4.5 Typical Reasoning Patterns
2-step pattern Reason-Refine Reason-Refine Reason-Debate
@@
3-step pattern Decompose-Refine-Reason Reason-Refine-Debate Reason-Decompose-Debate
@@
From the 2-step patterns, we observe that the Reason-Refine mode emerges frequently. Particularly in MATH and GPQA, which require massive mathematical calculations, this pattern compensates for the LLMs’ relatively poor calculation abilities, leading to more reliable results. In the 3-step patterns, operations like Decompose and Debate are frequently employed, which help break down challenging problems or facilitate discussions to explore potential solutions. Additionally, the Refine operation is often used before and after the Decompose and Debate steps, ensuring correct integration with preceding and succeeding reasoning processes. These typical reasoning patterns exhibit strong interpretability, further validating the capability of RLoT to generate task-specific logical structures that enhance LLM reasoning. For better understanding, we provide an example in Appendix B to illustrate how these logical structures empower the correct solution to a problem.
@@ 4.6 Ablation Studies, Analyses, and Discussions
To better justify our designs, we conducted a series of ablation studies, analyses, and discussions. We discuss the reliability of the self-evaluation state design in Appendix F. We analyze the necessity of employing RL to train the navigator model in Appendix G. We illustrate the role of the process reward model (PRM) in training the navigator in Appendix H. Also, we verify the effectiveness of our logic block designs and quantify the contribution of each logic block in Appendix I.
@@ 6 Conclusions
Appendix A Learning Curves
@@
Figure 2: Learning curves during RL training of all navigator models.
@@ Appendix C Computational Overhead and Latency
DeepSeek-R1-
Distill-Qwen-7B
Direct QA 879 11793 597 4686 541 4960 533 1364 808 5462 521 2094 436 1268 616 4518
@@ E.1 Test-time Scaling Baselines
Table 9: Extended baseline comparisons. The bold numbers indicate the best performance in each group of experiments, and the underlined numbers indicate the best baseline method.
LLM Method GSM8K GPQA StrategyQA Average
Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 63.89
DeAR (Xue et al., 2024) 88.38 36.16 71.13 65.22
Self-Refine (Madaan et al., 2023) 88.55 32.59 73.94 65.03
Tree search One-step-greedy 89.08 32.59 73.07 64.91
Basic MCTS 90.07 34.82 75.69 66.86
Litesearch (Wang et al., 2024a) 89.76 33.04 75.84 66.21
Q* (Wang et al., 2024b) 91.51 34.60 77.00 67.70
rStar (Qi et al., 2024) 91.96 38.62 77.15 69.24
AFlow (Zhang et al., 2024b) 91.36 37.72 76.56 68.55
DSPy) (Khattab et al., 2023) 88.32 34.15 74.09 65.52
Graph of Thoughts (GoT) (Besta et al., 2024) 88.96 33.04 75.42 65.81
Buffer of Thoughts (BoT) (Yang et al., 2024d) 92.35 34.51 74.33 67.06
RLoT (ours) 92.87 44.64 79.04 72.18
@@
Table 10: Token consumption and performance scores per question for representative baselines.
@@
Score Token Score Token Score Token Score Token
Qwen2.5-Instruct-7B DirectQA 91.58 240 31.25 705 68.85 301 63.89 415
Self-Refine 88.55 678 32.59 1784 73.94 939 65.03 1134
rStar 91.96 2636 38.62 4127 77.15 2987 69.24 3483
RLoT (ours) 92.87 1759 44.64 3001 79.04 2341 72.18 2367
@@ E.2 Math Domain Baselines
Appendix F Analyses on the Reliability of Self-evaluation State
@@
Mark has a garden with
flowers. He planted plants
of three different colors in it.
Ten of them are yellow,
and there are 80% more of
those in purple. There are only
25% as many green flowers as
there are yellow and purple
flowers. How many flowers
does Mark have in his garden?
@@ Appendix G Analyses on the Role of RL in Training the Navigator
Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 66.62
Repeated refine 87.72 32.37 72.34 66.65
Repeated debate 89.99 33.04 72.05 67.51
Fixed logic sequence 88.38 36.16 71.13 67.47
Supervise-trained navigator 89.84 36.83 70.31 67.94
LLM as navigator 89.16 37.95 74.38 69.17
RL-trained navigator (ours) 92.87 44.64 79.04 73.34
@@
Appendix H Analyses on the Role of Process Reward Model (PRM)
As literature (Zhang et al., 2025) indicates that math-trained PRMs perform well on MMLU-STEM, and the performance of PRMs in math and STEM is generally correlated. Our navigator, which is trained with a math PRM, achieves strong performance on GPQA (STEM). This also reflects the generalization capability of PRM across domains. We also test an ORM to train the navigator.
@@ Appendix H Analyses on the Role of Process Reward Model (PRM)
Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 66.62
RLoT with ORM 91.96 41.52 73.94 70.71
RLoT with PRM (ours) 92.87 44.64 79.04 73.34

### v3

@@ Abstract
Despite rapid advancements in large language models (LLMs), the token-level autoregressive nature constrains their complex reasoning capabilities. To enhance LLM reasoning, inference-time techniques, including Chain/Tree/Graph-of-Thought(s), successfully improve the performance, as they are fairly cost-effective by guiding reasoning through external logical structures without modifying LLMs’ parameters. However, these manually predefined, task-agnostic frameworks are applied uniformly across diverse tasks, lacking adaptability. To improve this, we propose RL-of-Thoughts (RLoT), where we train a lightweight navigator model with reinforcement learning (RL) to generate task-adaptive logical structures at inference time, enhancing LLM reasoning. Specifically, we design five basic logic blocks from the perspective of human cognition. During the reasoning process, the trained RL navigator dynamically selects the suitable logic blocks and combines them into task-specific logical structures according to problem characteristics. Experiments across multiple reasoning benchmarks (AIME, MATH, GPQA, etc.) with multiple LLMs (GPT, Llama, Qwen, and DeepSeek) illustrate that RLoT outperforms established inference-time techniques in most cases and improves up to 13.4% in challenging situations. Remarkably, with less than 3K parameters, our RL navigator is able to make sub-10B LLMs comparable to 100B-scale counterparts. Moreover, the RL navigator demonstrates strong transferability: a model trained on one specific LLM-task pair can effectively generalize to unseen LLMs and tasks. Our code is open-source at https://github.com/tsinghua-fib-lab/RL-LLM-Reasoning.
@@ 3.2 LLM Reasoning as MDP
A2: Clarity for further reasoning B2: Alternative methods in further reasoning C2: Completeness within the step
@@ 3.3 Training of the Navigator Model
Within the MDP framework outlined above, given a specific LLM type and a kind of reasoning task, we train the navigator model. Under this formulation, the training process constitutes a standard RL problem within a discrete action space. Consequently, our framework is algorithm-agnostic, allowing for the employment of arbitrary off-the-shelf RL algorithms for training. To enhance the learning for challenging reasoning tasks, we extract hard questions from the training set of the target task, i.e., questions that the LLM cannot answer when directly prompted. Then, we use these problems for training the navigator model, from which we randomly select one in each episode and repeat it multiple times.
@@ 4.1 Experimental Settings
4.2 Training Details
In our implementation, the navigator model is a simple three-layer multilayer perceptron (MLP) with the Dueling Network architecture (Wang et al., 2016). The model merely contains 2,566 parameters in total, where the lightweight design ensures efficient training and inference. We employ the Double-Dueling-DQN algorithm (Mnih et al., 2015; Van Hasselt et al., 2016; Wang et al., 2016) for optimization. By integrating Double Q-learning to mitigate value overestimation and the Dueling architecture to separate state and advantage representations, this algorithm significantly improves stability during the training process. We train the navigator model for 3,000 episodes with Qwen2.5-14B-Instruct on the MATH benchmark, where the learning curves presented in Figure 2 and Appendix A.1 indicate strong convergence. Furthermore, in Appendix A.2, we investigate the impact of employing alternative RL algorithms and extending training episodes, where the results consistently demonstrate similar convergence.
Figure 2: Learning curves during RL training of navigator models.
@@ 4.2 Training Details
DeepSeek-R1- Distill-Qwen-7B Direct QA 46.67 60.00 92.27 95.74 54.47 83.61 79.48 73.18
@@ 4.3 Overall Performance
We use the obtained navigator model to enhance the reasoning of multiple LLMs across different reasoning tasks. Following existing works like CoT-SC, we perform multiple repeated inferences for each task and filter out the trails that do not meet self-consistency, enhancing the robustness of reasoning, where the results are presented in Table 2. For the baseline performance, we prioritize using the results reported in the official technical reports of each LLM (Yang et al., 2024a; Dubey et al., 2024; Hurst et al., 2024), if available. Otherwise, we evaluate the performance through our own experiments following the settings established in previous works (Fu et al., 2023; Yang et al., 2024b; Yao et al., 2023).
@@ 4.6 Typical Reasoning Patterns
Instance number 5000 448 687
Major 2-step pattern Reason-Refine Reason-Refine Reason-Debate
@@
Major 3-step pattern Decompose-Refine-Reason Reason-Refine-Debate Reason-Decompose-Debate
@@
From the 2-step patterns, we observe that the Reason-Refine mode emerges frequently. Particularly in MATH and GPQA, which require massive mathematical calculations, this pattern compensates for the LLMs’ relatively poor calculation abilities, leading to more reliable results. In the 3-step patterns, operations like Decompose and Debate are frequently employed, which help break down challenging problems or facilitate discussions to explore potential solutions. Additionally, the Refine operation is often used before and after the Decompose and Debate steps, ensuring correct integration with preceding and succeeding reasoning processes. These typical reasoning patterns exhibit strong interpretability, further validating the capability of RLoT to generate task-specific logical structures that enhance LLM reasoning. For better understanding, we provide an example in Appendix B.1 to illustrate how these logical structures empower the correct solution to a problem.
@@ 4.7 Ablation Studies, Analyses, and Discussions
To better justify our designs, we verify the effectiveness of our logic block designs with ablation studies. To examine the impact of each block, we remove each of them and train a navigator, where results in Table 6 confirm that all blocks are effective.
Table 6: Ablation study results for each logic block.
LLM Ablation MATH GPQA StrategyQA Average
Qwen2.5-Instruct-7B w/o Decompose 75.42 31.92 77.00 61.45
w/o Debate 74.02 36.61 77.58 62.74
w/o Refine 75.76 41.29 72.93 63.33
Full RLoT 76.70 44.64 79.04 66.79
GPT-4o-mini w/o Decompose 76.08 41.74 79.33 65.72
w/o Debate 74.68 45.31 78.75 66.25
w/o Refine 76.44 51.56 74.38 67.46
Full RLoT 77.36 54.02 82.68 71.35
Also, we conducted a series of analyses, and discussions. We quantify the contribution of each logic block in Appendix I. We discuss the reliability of the self-evaluation state design in Appendix F. We analyze the necessity of employing RL to train the navigator model in Appendix G. We illustrate the role of the process reward model (PRM) in training the navigator in Appendix H.
@@ 6 Conclusions
Appendix A Training Details
A.1 Learning Curves
@@
Figure 3: Learning curves during RL training of navigator models.
A.2 Analysis
In Figure 4, we extend the training duration to 5,000 episodes. We observe that the model essentially reaches convergence around 3,000 episodes, with only marginal performance gains achieved thereafter. Additionally, we explore training the navigator model using the Double-DQN algorithm. While the training stability decreases slightly due to the absence of the Dueling network architecture, the model ultimately converges to a performance level comparable to that of the Double-Dueling-DQN.
Figure 4: 5k episodes learning curves during RL training with different algorithms.
@@ B.1 Correct Cases
B.2 Failure Cases
To provide a comprehensive understanding of the boundaries of RLoT, we also analyze typical failure cases observed during our experiments. Common failures can be categorized into two primary modes: (1) Navigator input errors caused by self-evaluation noise, and (2) Navigator output errors caused by suboptimal action selection.
As shown in Figure 6, failure happens due to self-evaluation noise. The robustness of the navigator depends partly on the accuracy of the state representation derived from the LLM self-evaluation. In this case, the LLM generates an incorrect equation (wrongly deriving 3q=24p363q=24p-36) but fails to identify this error during the self-evaluation phase, hallucinating a "True" status for correctness of modeling. This noise in the state input misleads the navigator: instead of selecting the necessary "Refine" action to correct the mistake, the Navigator perceives the current step as valid and selects "Debate" to explore alternative plans.
Figure 6: A failure case study that is caused by wrong state self-evaluation.
As shown in Figure 7, failure happens due to action selection error, where the navigator selects a suboptimal action despite receiving accurate state information. In this case, the LLM derives a correct intermediate result ($\cos(2\pi/7)+\dots=-1/2$) and correctly evaluates it as "True". However, the navigator triggers a "Refine" action instead of proceeding with reasoning or termination. This unnecessary intervention forces the LLM to modify a correct step, leading to "over-correction" where valid reasoning is replaced with an erroneous identity, ultimately causing the solution to fail.
Figure 7: A failure case study that is caused by wrong action selection.
While these failure modes exist, our extensive quantitative results demonstrate that RLoT achieves performance gains over baselines, indicating that these failure cases represent a minority of the inference trajectories. In our design, the RL training process effectively optimizes the navigator to minimize action selection errors, and the structured self-evaluation prompt is designed to mitigate state noise errors as much as possible.
@@ Appendix C Computational Overhead and Latency
DeepSeek-R1- Distill-Qwen-7B Direct QA 879 11793 597 4686 541 4960 533 1364 808 5462 521 2094 436 1268 616 4518
@@ E.1 Test-time Scaling Baselines
  • Least-to-Most (Zhou et al., 2023): Ask LLMs to break down a complex problem into a series of simpler subproblems and then solve them in sequence.
  • SelfCheck (Miao et al., 2024): Utilize LLMs to check their own outputs and use the results of these checks to improve question-answering performance by conducting weighted voting on multiple solutions to the question. Here we adopt voting with 4 candidates.
@@
Table 10: Extended baseline comparisons. The bold numbers indicate the best performance in each group of experiments, and the underlined numbers indicate the best baseline method. All results are averaged across 5 repeated runs with standard deviation.
LLM Method MATH-500 GSM8K GPQA StrategyQA Average
Qwen2.5-Instruct-7B DirectQA 74.96±1.37 90.95±1.06 31.16±0.23 68.53±0.53 66.40±0.46
Self-Refine 72.32±1.20 88.39±0.38 32.23±0.61 74.06±0.79 66.75±0.33
Least-to-Most 73.28±0.84 89.10±0.82 31.79±0.54 73.60±0.64 66.94±0.25
SelfCheck 74.36±1.60 90.24±1.09 33.84±0.61 72.87±0.77 67.83±0.35
DeAR 69.92±0.88 87.84±0.54 35.71±0.37 71.03±0.85 66.13±0.39
Tree search One-step-greedy 71.24±1.43 89.49±0.99 32.63±0.50 72.43±0.56 66.45±0.54
Basic MCTS 72.40±0.79 89.51±0.72 35.13±0.59 75.72±0.80 68.19±0.19
Litesearch 72.52±1.29 89.86±0.86 33.17±0.30 76.04±0.65 67.90±0.52
Q* 73.76±0.81 92.28±0.61 34.38±0.76 76.80±0.99 69.30±0.50
rStar 75.88±1.02 92.05±0.89 38.17±0.32 77.09±0.83 70.80±0.68
AFlow 75.72±1.16 90.95±0.98 37.37±0.63 77.09±0.56 70.28±0.13
DSPy 72.48±0.86 88.75±0.77 34.11±0.68 73.89±0.94 67.31±0.49
Graph of Thoughts (GoT) 73.24±1.06 88.76±1.07 33.26±0.66 75.69±0.91 67.74±0.39
Buffer of Thoughts (BoT) 75.16±0.74 91.98±0.64 34.91±0.58 74.38±0.44 69.11±0.32
RLoT (ours) 76.88±1.35 92.95±0.63 44.78±0.71 79.56±0.64 73.54±0.18
@@
Table 11: Token consumption, solving time (minutes) per question, and performance scores for representative baselines.
@@
Score Token Time Score Token Time Score Token Time Score Token Time
Qwen2.5-Instruct-7B DirectQA 91.58 240 0.05 31.25 705 0.17 68.85 301 0.05 63.89 415 0.09
Self-Refine 88.55 678 0.12 32.59 1784 0.41 73.94 939 0.16 65.03 1134 0.23
Litesearch 89.76 912 0.20 33.04 2139 0.55 75.84 1228 0.24 66.21 1426 0.33
rStar 91.96 2636 0.77 38.62 4127 1.44 77.15 2987 0.82 69.24 3483 1.01
RLoT (ours) 92.87 1759 0.61 44.64 3001 0.95 79.04 2341 0.53 72.18 2367 0.70
@@ E.2 Math Domain Baselines
Appendix F Analyses on the Self-evaluation State
F.1 Reliability
@@
Mark has a garden with flowers. He planted plants of three different colors in it. Ten of them are yellow, and there are 80% more of those in purple. There are only 25% as many green flowers as there are yellow and purple flowers. How many flowers does Mark have in his garden?
@@
F.2 Noise Impact
To investigate the navigator’s reliance on accurate state perceptions, we introduced synthetic noise into the self-evaluation state vectors during inference. As shown in Table 14, replacing the state with random values results in a significant performance drop, falling even below DirectQA. This confirms that the navigator actively utilizes the state information for decision-making rather than following a fixed, state-agnostic policy. However, the framework exhibits commendable robustness. Under extreme conditions with 50% noise (where half of the state bits are flipped), RLoT still maintains an average accuracy of 64.99%, being comparable to the DirectQA baseline. As the noise decreases to 30%, performance recovers rapidly to 69.11%.
Table 14: Impact of noise in self-evaluation states.
LLM State GSM8K GPQA StrategyQA Average
Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 63.89
Random 89.76 28.13 65.50 61.13
50% noise 91.74 34.82 68.41 64.99
30% noise 92.27 43.75 71.32 69.11
Normal 92.87 44.64 79.04 72.18
Meanwhile, to minimize potential noise inherent in self-evaluation, we deliberately designed the evaluation criteria as a binary classification task rather than a multi-grade continuous scoring. Continuous or fine-grained scoring (e.g., 1-10 scales) requires precise calibration, which is challenging for LLMs. In contrast, a binary standard simplifies the decision boundary, significantly reducing the cognitive load on the LLM. This design choice makes the evaluation process more robust, ensuring that the state vector remains a stable and reliable signal for the navigator even with less capable LLMs.
F.3 Alternative Designs
@@ Appendix G Analyses on the Role of RL in Training the Navigator
Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 63.89
Repeated refine 87.72 32.37 72.34 64.14
Repeated debate 89.99 33.04 72.05 65.03
Fixed logic sequence 88.38 36.16 71.13 65.22
Supervise-trained navigator 89.84 36.83 70.31 65.66
LLM as navigator 89.16 37.95 74.38 67.16
RL-trained navigator (ours) 92.87 44.64 79.04 72.18
@@
Appendix H Analyses on the Process Reward Model (PRM)
H.1 The Role of PRM
To investigation the role of PRM, we also test an ORM to train the navigator.
@@ H.1 The Role of PRM
Qwen2.5-Instruct-7B DirectQA 91.58 31.25 68.85 63.89
RLoT with ORM 91.96 41.52 73.94 69.14
RLoT with PRM (ours) 92.87 44.64 79.04 72.18
@@
H.2 Generalization Capability of PRM
Recent studies suggest that process reward models (PRMs) trained on mathematical data exhibit strong transferability to related domains such as STEM (Zhang et al., 2025). To empirically verify this and justify our reward design, we conducted a PRM calibration analysis. We calculated the correlation between the intermediate PRM scores assigned at various reasoning steps and the correctness of the final answer.
Table 18: PRM calibration.
LLM Correlation with Correctness MATH GPQA StrategyQA
Qwen2.5-Instruct-7B Step1 PRM 0.199 0.184 0.157
Step2 PRM 0.222 0.213 0.195
Step3 PRM 0.236 0.215 0.202
Step4 PRM 0.240 0.228 0.218
Average 0.224 0.210 0.193
As presented in Table 18, the results reveal two insights. First, there is a consistent positive correlation between intermediate PRM scores and the final outcome across all tested benchmarks. This confirms that the PRM effectively gauges the quality of intermediate steps, validating its suitability as a dense reward signal for training the navigator. Second, although the PRM was trained primarily on mathematical data, its scores exhibit a notable correlation with solution correctness on out-of-domain tasks, including GPQA (STEM) and StrategyQA (commonsense). This demonstrates the PRM’s intrinsic ability to generalize its verification logic across different domains.
This calibration analysis explains the underlying mechanism of our method’s transferability. Because the math-trained PRM provides reliable feedback even on unseen tasks like GPQA and StrategyQA, the navigator trained with these rewards can successfully learn effective reasoning policies across diverse domains.
H.3 Impact of PRM Quality